OpenStudy (anonymous):

Solve for x: 2 over 3 (x - 4) = 2x. 2 -2 -8 -4 Does anyone know this question, i need help

OpenStudy (leader):

\[\frac{ 2 }{ 3}(x-4)=2x\]

OpenStudy (leader):

x-4=3x

OpenStudy (leader):

2x=-4

OpenStudy (leader):

x=-2

OpenStudy (leader):

@DieselLife

OpenStudy (leader):

Fan and medal?

OpenStudy (anonymous):

how?

OpenStudy (leader):

there is a best response button -------------> And put your mouse over my name and let it load and it will say become a fan :)

OpenStudy (anonymous):

ok and how for the medal?

OpenStudy (anonymous):

and thanks for the answer @Leader